OpenStudy (anonymous):

b+c=10/a solve for a

OpenStudy (anonymous):

a=10/(b+c)

OpenStudy (anonymous):

\[b+c = \frac{10}{a}\]\[a(b+c) = 10\]\[a = \frac{10}{(b+c)}\]

OpenStudy (anonymous):

\[a=\frac{10}{b+c} \]

OpenStudy (anonymous):

can u guys do all the steps?

OpenStudy (anonymous):